# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 4.4. Mickelsson algebras and difference operators
Recall that $\mathbf{u}_{i_{1}}^{-1}\ldots\mathbf{u}_{i_{k}}^{-1}$ is proportional to $Pv_{i_{1}}P\ldots Pv_{i_{k}}P$. By ˜3.9, the action of a monomial on the latter is
$Y_{1}^{j_{1}}\ldots Y_{k}^{j_{k}}\cdot Pv_{i_{1}}P\ldots Pv_{i_{k}}P=P\Omega^{j_{1}}(v_{i_{1}})P\ldots P\Omega^{j_{k}}(v_{i_{k}})P$
(it is well-defined once the function is symmetric). The result follows by ˜4.6. ∎
5. Kostant-Whittaker reduction and Yangians
In this section, we will study a variant of the constructions of Section˜4. Our main goal is to relate the shifted Yangian of $\mathfrak{sl}_{2}$ to the Whittaker analog of the Mickelsson algebra for $S^{\bullet}\mathbf{C}^{N}$.
@@ 5. Kostant-Whittaker reduction and Yangians
(5.1) $\psi\colon\mathfrak{n}_{+}\rightarrow\mathbf{C},\qquad\psi(E_{ij})=\delta_{i+1,j}\psi$
@@
$\mathfrak{n}_{+}^{\psi}:=\{x-\psi(x)|x\in\mathfrak{n}_{+}\}\subset\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}).$
For any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$, denote by $X/\mathfrak{n}_{+}^{\psi}$ the quotient by the right action of $\mathfrak{n}_{+}^{\psi}$. Recall that by Kostant’s construction [Kos78], the space of invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\!\!/_{\psi}\,N_{+}:=(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\mathfrak{n}_{+}^{\psi})^{N_{+}}$ is isomorphic to the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. It admits the following categorification.
Definition 5.1.
[BF08] The Kostant-Whittaker reduction is a functor
$\mathrm{res}^{\psi}_{+}\colon\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})\rightarrow{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}},\qquad X\mapsto X/\!\!/_{\psi}\,N_{+}:=(X/\mathfrak{n}_{+}^{\psi})^{N_{+}}.$
There is a Whittaker analog of the extremal projector (4.6) introduced in [Kal25] under the name of the Kirillov projector.
@@
$\begin{split}(x-\psi(x))P_{\psi}&=0,\ x\in\mathfrak{n}_{+},\\ P_{\psi}(E_{ij}+\delta_{ij}\hbar(N-i))&=0,\ 1\leq j\leq i\leq N-1,\\ (P_{\psi})^{2}=P_{\psi},&\qquad P_{\psi}\Big{|}_{(X/\mathfrak{n}_{+}^{\psi})^{N_{+}}}=\mathrm{id}.\end{split}$
@@
In the notations of [Kal25, Definition 5.2], we can apply (1.6) to $P_{\mathfrak{m}_{N}}^{\psi}(N\hbar-\hbar,\ldots,\hbar)$; by [Kal25, Theorem 5.3], it satisfies the properties of the theorem for $\psi(E_{ij})=\delta_{i+1,j}$. For general $\psi\in\mathbf{C}^{*}$, we can apply the automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ defined by $x\mapsto x/\psi$ for $x\in\mathfrak{gl}_{N}$ and $\hbar\mapsto\hbar/\psi$. ∎
As in Section˜4, we use the following.
Notation 5.3.
@@
(5.2) $\mathfrak{b}^{\rho}_{\mathrm{mir}}\backslash X/\mathfrak{n}_{+}^{\psi}\xrightarrow{\sim}X/\!\!/_{\psi}\,N_{+},\qquad[x]\mapsto P_{\psi}xP_{\psi},$
@@
$\mathfrak{b}^{\rho}_{\mathrm{mir}}:=\mathrm{span}\left(E_{ij}+\delta_{ij}\cdot\hbar(N-i)|1\leq j\leq i\leq N-1\right)$
@@
(5.3) $\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes V\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/_{\psi}\,N_{\pm},\qquad v\mapsto P_{\psi}vP_{\psi}.$
We can compute explicitly this trivialization for the vector representation $\mathbf{C}^{N}$ using the canonical homomorphism (3.3). The following can be proved analogously to [Kal25, Proposition 6.18].
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}).$
@@
Combining (5.3), ˜5.4 and Proposition˜3.7, we conclude that the map
@@
is an isomorphism of modules over $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]$. By Proposition˜4.7, we have
@@
By expanding in powers of $u$ as in Remark˜4.8, we can conclude that an endomorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ commuting with the left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action commutes with the right $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action if and only if it commutes with $\Omega$. Therefore, such an endomorphism commutes with the algebra action of $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))$, giving a canonical isomorphism
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}).$
5.1. Kostant-Mickelsson-Whittaker algebras
Similarly to (4.12), the Kostant-Whittaker reduction is endowed with a monoidal structure via
(5.4) $X/\!\!/_{\psi}\,N_{+}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}Y/\!\!/_{\psi}\,N_{+}\xrightarrow{\sim}(X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}Y)/\!\!/_{\psi}\,N_{+},\qquad[x]\otimes[y]\mapsto[x\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}y],$
see [BF08]. In particular, we have an analog of ˜4.10.
@@ 5.1. Kostant-Mickelsson-Whittaker algebras
Let $A$ be an algebra object in $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$. The Kostant-Mickelsson-Whittaker algebra associated to $A$ is the $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodule $A/\!\!/_{\psi}\,N_{+}$ with the product
$A/\!\!/_{\psi}\,N_{+}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}A/\!\!/_{\psi}\,N_{+}\xrightarrow{\sim}(A\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}A)/\!\!/_{\psi}\,N_{+}\rightarrow A/\!\!/_{\psi}\,N_{+},$
@@
$P_{\psi}aP_{\psi}\otimes P_{\psi}bP_{\psi}\mapsto P_{\psi}aP_{\psi}bP_{\psi},\ a,b\in A.$
We will explicitly compute the Kostant-Mickelsson-Whittaker algebras $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes T(\mathbf{C}^{N})/\!\!/_{\psi}\,N_{+}$, where $T(\mathbf{C}^{N}):=\bigoplus_{i=0}^{\infty}(\mathbf{C}^{N})^{\otimes i}$ is the tensor algebra, and $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$.
Recall the action of the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ from Section˜3.1. It descends to an action on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}$. Recall the central elements $A(u)=\sum_{i}A_{i}u^{N-i}$ of (1.11). Define the elements $A^{(j)}_{i}\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]$ for $1\leq j\leq k+1$ inductively via $A^{(1)}_{i}=A_{i}$ for all $i$ and
$A_{i}^{(j+1)}-\Omega_{j}A_{i-1}^{(j+1)}=A_{i}^{(j)}-(\Omega_{j}-\hbar)A_{i-1}^{(j)},\qquad A_{i}^{(j)}=\delta_{i0},\ i\leq 0.$
Using Remark˜4.8, it can be rewritten as
(5.5) $A^{(j+1)}(u)=\frac{u-\Omega_{j}+\hbar}{u-\Omega_{j}}A^{(j)}(u),\qquad A^{(j)}(u):=\sum_{i=0}^{N}A^{(j)}_{i}u^{N-i}.$
@@
(5.6) $A^{(j)}(u)=A(u)\prod_{a=1}^{j-1}\frac{u-\Omega_{a}+\hbar}{u-\Omega_{a}}.$
Consider the algebra $\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]$ with product
(5.7) $f(\Omega_{1},\ldots,\Omega_{k})\otimes A(u)$ $\mapsto A^{(k+1)}(u)f(\Omega_{1},\ldots,\Omega_{k}),$
(5.8) $f(\Omega_{1},\ldots,\Omega_{k})\otimes g(\Omega_{1},\ldots,\Omega_{k+l})$ $\mapsto f(\Omega_{1},\ldots,\Omega_{k})g(\Omega_{k+1},\ldots,\Omega_{k+l}),$
for $f(\Omega_{1},\ldots,\Omega_{k})\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]$ and $g(\Omega_{1},\ldots,\Omega_{l})\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{l}]$ regarded as elements in the corresponding graded components.
The following is a generalization of Proposition˜5.5.
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1\ldots k)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1\ldots k)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{k=0}^{\infty}(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\mathrm{End}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}^{\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}}(\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes(\mathbf{C}^{N})^{\otimes k})$
    the space of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-linear endomorphism of the free left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module $\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes(\mathbf{C}^{N})^{\otimes k}$ commuting with the action of the spherical subalgebra $\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$ of $\mathfrak{H}^{\hbar}_{k}$. Then we have an isomorphism
    $\mathrm{End}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}^{\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}}(\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes(\mathbf{C}^{N})^{\otimes k})\cong\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}).$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}\cong(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+})\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\ldots\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+})\cong$
$\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1}]/(A(\Omega_{1}))\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\ldots\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{k}]/(A(\Omega_{k})).$
By Proposition˜4.7, the right action of the center on each tensor component satisfies
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]/(A^{(j)}(\Omega_{j})|j=1,\ldots,k)$
@@
By ˜A.2, they generate the ring of symmetric functions in $\{\Omega_{i}\}$. Therefore, we have a desired isomorphism. ∎

### v4

@@ 5.4 Mickelsson algebras and difference operators
Recall that $\mathbf{u}_{i_{1}}^{-1}\cdots\mathbf{u}_{i_{k}}^{-1}$ is proportional to $Pv_{i_{1}}P\cdots Pv_{i_{k}}P$. By Corollary 4.9, the action of a monomial on the latter is
$Y_{1}^{j_{1}}\cdots Y_{k}^{j_{k}}\cdot Pv_{i_{1}}P\cdots Pv_{i_{k}}P=P\Omega^{j_{1}}(v_{i_{1}})P\cdots P\Omega^{j_{k}}(v_{i_{k}})P$
(it is well-defined once the function is symmetric). The result follows by Lemma 5.6. ∎
6 Kostant–Whittaker reduction and Yangians
In this section, we will study the Whittaker variant of the constructions from Section 5. We use the action of degenerate affine Hecke algebras from Section 4 to find a presentation of the Kostant–Whittaker reduction of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ as a certain quotient of symmetric polynomial in $k$ variables over the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$, see Corollary 6.8. It is used to construct a homomorphism from the non-negative part $\mathrm{Y}_{-N,N}(\mathfrak{sl}_{2})^{\geq 0}$ of the shifted Yangian to the Kostant–Whittaker reduction of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ by using the results of Section 3. The main construction of the paper, Theorem 8.10, is essentially a combination of the one from this section with the corresponding dual counterpart of Appendix B. In Section 6.3, we discuss a mirabolic analog of the results from previous subsections.
@@ 6 Kostant–Whittaker reduction and Yangians
$\psi\colon\ \mathfrak{n}_{+}\rightarrow\mathbf{C},\qquad\psi(E_{ij})=\delta_{i+1,j}\psi$ (6.1)
@@
$\mathfrak{n}_{+}^{\psi}:=\{x-\psi(x)\mid x\in\mathfrak{n}_{+}\}\subset\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}).$
For any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$, denote by $X/\mathfrak{n}_{+}^{\psi}$ the quotient by the right action of $\mathfrak{n}_{+}^{\psi}$. Recall that by Kostant’s construction [33], the space of invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\!\!/_{\psi}\,N_{+}:=\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\mathfrak{n}_{+}^{\psi}\bigr)^{N_{+}}$ is isomorphic to the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. It admits the following categorification.
Definition 6.1 ([3]).
The Kostant–Whittaker reduction is a functor
$\mathrm{res}^{\psi}_{+}\colon\ \mathrm{HC}_{\hbar}(\mathrm{GL}_{N})\rightarrow{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}},\qquad X\mapsto X/\!\!/_{\psi}\,N_{+}:=\bigl(X/\mathfrak{n}_{+}^{\psi}\bigr)^{N_{+}}.$
There is a Whittaker analog of the extremal projector (5.5) introduced in [27] under the name of the Kirillov projector.
@@
$\begin{split}&(x-\psi(x))P_{\psi}=0,\qquad x\in\mathfrak{n}_{+},\\ &P_{\psi}(E_{ij}+\delta_{ij}\hbar(N-i))=0,\qquad 1\leq j\leq i\leq N-1,\\ &(P_{\psi})^{2}=P_{\psi},\qquad P_{\psi}|_{(X/\mathfrak{n}_{+}^{\psi})^{N_{+}}}=\mathrm{id}.\end{split}$
@@
In the notations of [27, Definition 5.2], we can apply (2.5) to $P_{\mathfrak{m}_{N}}^{\psi}(N\hbar-\hbar,\dots,\hbar)$; by [27, Theorem 5.3], it satisfies the properties of the theorem for $\psi(E_{ij})=\delta_{i+1,j}$. For general $\psi\in\mathbf{C}^{*}$, we can apply the automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ defined by $x\mapsto x/\psi$ for $x\in\mathfrak{gl}_{N}$ and $\hbar\mapsto\hbar/\psi$. ∎
As in Section 5, we use the following.
Notation 6.3.
@@
$\mathfrak{b}^{\rho}_{\mathrm{mir}}\backslash X/\mathfrak{n}_{+}^{\psi}\xrightarrow{\sim}X/\!\!/_{\psi}\,N_{+},\qquad[x]\mapsto P_{\psi}xP_{\psi},$ (6.2)
@@
$\mathfrak{b}^{\rho}_{\mathrm{mir}}:=\mathrm{span}(E_{ij}+\delta_{ij}\cdot\hbar(N-i)\mid 1\leq j\leq i\leq N-1).$
@@
$\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes V\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/_{\psi}\,N_{\pm},\qquad v\mapsto P_{\psi}vP_{\psi}.$ (6.3)
We can compute explicitly this trivialization for the vector representation $\mathbf{C}^{N}$ using the canonical homomorphism (4.2). The following can be proved analogously to [27, Proposition 6.18].
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr).$
@@
Combining (6.3), Lemma 6.4 and Proposition 4.7, we conclude that the map
@@
is an isomorphism of modules over $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]$. By Proposition 5.7, we have
@@
By expanding in powers of $u$ as in Remark 5.8, we can conclude that an endomorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ commuting with the left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action commutes with the right $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action if and only if it commutes with $\Omega$. Therefore, such an endomorphism commutes with the algebra action of $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))$, giving a canonical isomorphism
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr),$
as claimed. ∎
6.1 Kostant–Mickelsson–Whittaker algebras
Similarly to (5.11), the Kostant–Whittaker reduction is endowed with a monoidal structure via
$\begin{split}X/\!\!/_{\psi}\,N_{+}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}Y/\!\!/_{\psi}\,N_{+}&\xrightarrow{\sim}\bigl(X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}Y\bigr)/\!\!/_{\psi}\,N_{+},\\ [x]\otimes[y]&\mapsto\bigl[x\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}y\bigr],\end{split}$ (6.4)
see [3]. In particular, we have an analog of Definition 5.10.
@@ 6.1 Kostant–Mickelsson–Whittaker algebras
Let $A$ be an algebra object in $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$. The Kostant–Mickelsson–Whittaker algebra associated to $A$ is the $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodule $A/\!\!/_{\psi}\,N_{+}$ with the product
$A/\!\!/_{\psi}\,N_{+}\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}A/\!\!/_{\psi}\,N_{+}\xrightarrow{\sim}\bigl(A\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}A\bigr)/\!\!/_{\psi}\,N_{+}\rightarrow A/\!\!/_{\psi}\,N_{+},$
@@
$P_{\psi}aP_{\psi}\otimes P_{\psi}bP_{\psi}\mapsto P_{\psi}aP_{\psi}bP_{\psi},\qquad a,b\in A.$
We explicitly compute the Kostant–Mickelsson–Whittaker algebras $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes T\bigl(\mathbf{C}^{N}\bigr)/\!\!/_{\psi}\,N_{+}$, where $T\bigl(\mathbf{C}^{N}\bigr):=\bigoplus_{i=0}^{\infty}\bigl(\mathbf{C}^{N}\bigr)^{\otimes i}$ is the tensor algebra, and $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$.
Recall the action of the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ from Section 4.1. It descends to an action on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}$. Recall the central elements $A(u)=\sum_{i}A_{i}u^{N-i}$ of (2.8). Define the elements $A^{(j)}_{i}\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]$ for $1\leq j\leq k+1$ inductively via $A^{(1)}_{i}=A_{i}$ for all $i$ and
$A_{i}^{(j+1)}-\Omega_{j}A_{i-1}^{(j+1)}=A_{i}^{(j)}-(\Omega_{j}-\hbar)A_{i-1}^{(j)},\qquad A_{i}^{(j)}=\delta_{i0},\qquad i\leq 0.$
Using Remark 5.8, it can be rewritten as
$A^{(j+1)}(u)=\frac{u-\Omega_{j}+\hbar}{u-\Omega_{j}}A^{(j)}(u),\qquad A^{(j)}(u):=\sum_{i=0}^{N}A^{(j)}_{i}u^{N-i}.$ (6.5)
@@
$A^{(j)}(u)=A(u)\prod_{a=1}^{j-1}\frac{u-\Omega_{a}+\hbar}{u-\Omega_{a}}.$ (6.6)
Consider the algebra $\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]$ with product
$f(\Omega_{1},\dots,\Omega_{k})\otimes A(u)\mapsto A^{(k+1)}(u)f(\Omega_{1},\dots,\Omega_{k}),$ (6.7)
$f(\Omega_{1},\dots,\Omega_{k})\otimes g(\Omega_{1},\dots,\Omega_{k+l})\mapsto f(\Omega_{1},\dots,\Omega_{k})g(\Omega_{k+1},\dots,\Omega_{k+l})$ (6.8)
for $f(\Omega_{1},\dots,\Omega_{k})\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]$ and $g(\Omega_{1},\dots,\Omega_{l})\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{l}]$ regarded as elements in the corresponding graded components.
The following is a generalization of Proposition 6.5.
@@
    $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\bigoplus_{k=0}^{\infty}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigoplus_{k=0}^{\infty}\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+},$
@@
    $\mathrm{End}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}^{\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}}\bigl(\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\bigr)$
    the space of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-linear endomorphism of the free left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module $\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ commuting with the action of the spherical subalgebra $\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}$ of $\mathfrak{H}^{\hbar}_{k}$. Then we have an isomorphism
    $\mathrm{End}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}^{\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}}\bigl(\mathbf{Z}_{\mathfrak{gl}_{N}}\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\bigr)\cong\mathrm{End}_{{}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathrm{BMod}_{\mathbf{Z}_{\mathfrak{gl}_{N}}}}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}\bigr).$
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}$
$\qquad\cong\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr)\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\dots\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}\bigr)$
$\qquad\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1}]/(A(\Omega_{1}))\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\dots\otimes_{\mathbf{Z}_{\mathfrak{gl}_{N}}}\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{k}]/(A(\Omega_{k})).$
By Proposition 5.7, the right action of the center on each tensor component satisfies
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}/\!\!/_{\psi}\,N_{+}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\dots,\Omega_{k}]/\bigl(A^{(j)}(\Omega_{j})\mid j=1,\dots,k\bigr)$
@@
By Lemma A.2, they generate the ring of symmetric functions in $\{\Omega_{i}\}$. Therefore, we have a desired isomorphism. ∎